OpenStudy (love.kat123):

I beg you. If you have time, please help me with these algebra questions? medal+fan

OpenStudy (love.kat123):

Simplify the radical below. \[\sqrt[4]{32z ^{8}}\]

OpenStudy (love.kat123):

Rewrite the radicand as two factors, one of which is a perfect cube. \[\sqrt[3]{250}\]

OpenStudy (love.kat123):

both i dont know what do at allim using high lighters green for what i know yellow for im learning and red for the things i don know at all im mostly using red

OpenStudy (love.kat123):

okay

OpenStudy (love.kat123):

erh yea........... no

OpenStudy (love.kat123):

i guess idk im going to fail

OpenStudy (love.kat123):

:/

OpenStudy (loser66):

Ok, I restart by simple expression. Let your problem aside. Trust me, just 1 time, you get the concept and no more failure. Try or not??

OpenStudy (love.kat123):

doesn't matter im going to fail either way :/ but ill try

OpenStudy (loser66):

 I want you name the number inside and outside relate to 2. In this case inside is 5 out side is 2, got it? Tell me, \(\huge \sqrt[3 ] {2^8}\) which number is inside, which is out side. I am trying to make you familiar with the new concept. just answer me

OpenStudy (love.kat123):

okay

OpenStudy (loser66):

\[\huge \sqrt[4 ] {2^9}\] outside? inside?

OpenStudy (love.kat123):

4 and 9

OpenStudy (love.kat123):

3 and 7

OpenStudy (love.kat123):

3 and 8

OpenStudy (love.kat123):

..........................................

OpenStudy (love.kat123):

confusing

OpenStudy (love.kat123):

girl

OpenStudy (loser66):

how old are you?

OpenStudy (love.kat123):

1.3

OpenStudy (love.kat123):

XD 15